# Entropy-Triggered Retraining as Nonequilibrium Entropy Production in Deployed Machine Learning Systems

### v1

@@ Abstract
Machine learning models deployed in nonstationary environments inevitably experience performance degradation due to data drift. While numerous drift detection heuristics exist, most lack a dynamical interpretation and offer limited guidance on how retraining should be balanced against operational cost. In this work, we propose an entropy-based retraining framework grounded in nonequilibrium statistical physics. Modeling drift as probability flow governed by a Fokker–Planck equation, we quantify model–data mismatch using a time-evolving Kullback–Leibler divergence. We show that its rate of change admits an entropy-balance decomposition featuring a nonnegative entropy production term driven by probability currents. Guided by this theory, we evaluate entropy-triggered retraining in a controlled nonstationary classification setting and demonstrate near-maximum-frequency baseline predictive performance while reducing retraining frequency by an order of magnitude relative to daily and label-based strategies.
@@ 1 Introduction
  • We empirically demonstrate that entropy-triggered retraining achieves near-maximum-frequency baseline predictive performance while reducing retraining frequency by an order of magnitude relative to daily and label-based retraining strategies.
@@
Section 2 introduces a probability-flow model of drift. Section 3 defines mismatch entropy and derives its evolution along the flow. Section 4 interprets retraining as an external control that resets mismatch. Section 5 evaluates entropy-triggered retraining experimentally. Appendix A provides the entropy-balance derivation.
@@ 2 Drift as Probability Flow
Let $X_{t}\in\mathbb{R}^{d}$ denote the feature vector observed at deployment time $t$. We model its evolution by the stochastic differential equation
@@
where $a$ is a drift field, $B$ a diffusion matrix, and $W_{t}$ standard Brownian motion.
Under standard regularity assumptions, the probability density $p(x,t)$ of $X_{t}$ satisfies the Fokker–Planck equation
@@
The Fokker–Planck equation provides a canonical description of the evolution of probability densities under stochastic dynamics and admits a rich analytical structure; see [Risken1989] for a comprehensive treatment.
Nonzero probability currents correspond to nonequilibrium dynamics and provide a mechanism by which a fixed deployed model becomes misaligned with its environment. From the perspective of nonequilibrium statistical physics, such currents are fundamental drivers of irreversibility and entropy production, encoding sustained deviations from equilibrium through persistent probability flow [Schnakenberg1976].
3 Mismatch Entropy as a Lyapunov Functional
@@
which measures the statistical divergence between the evolving deployment-time data distribution $p(x,t)$ and the fixed reference distribution associated with the deployment environment. From an information-theoretic perspective, relative entropy provides a natural and asymmetric measure of distributional mismatch that is independent of labels or model outputs [CoverThomas2006].
Differentiating $D(t)$ along the Fokker–Planck flow yields
@@
In reversible settings where the reference density aligns with a stationary distribution of the flow, relative entropy acts as a Lyapunov functional and dissipates along the dynamics. More generally, along nonequilibrium probability flows the entropy rate admits a decomposition analogous to the entropy balance relation of stochastic thermodynamics, featuring a nonnegative entropy production term driven by probability currents and a residual “housekeeping” term that captures sustained driving. Such entropy balance relations play a central role in stochastic thermodynamics for diffusion processes [Seifert2005], and closely related variational structures for Fokker–Planck dynamics are developed in [JordanKinderlehrerOtto1998]. The full derivation of the decomposition used here is provided in Appendix A.
@@
When $q_{\mathrm{ref}}$ coincides with an appropriate stationary reference for the flow, the mismatch entropy $D(t)$ behaves as a Lyapunov functional and decays monotonically. Under driven nonequilibrium drift, $D(t)$ need not be monotone; nevertheless, its evolution is constrained by an entropy-balance relation with a nonnegative entropy production term, and increases in mismatch reflect the effect of sustained probability currents relative to the fixed deployed reference distribution.
@@
Figure 1: Evolution of mismatch entropy under nonstationary drift. The threshold defines entropy-triggered retraining events. Small negative empirical values arise from finite-sample estimation noise; the population Kullback–Leibler divergence is nonnegative.
@@
Sections 2–3 provide a simple implication: if the deployment-time distribution evolves under probability currents, then mismatch relative to a fixed deployed reference will generally change over time, and sustained drift produces nonequilibrium entropy production. Maintaining predictive performance therefore requires an intervention that updates the deployed reference. In deployed learning systems, this intervention is retraining.
4 Retraining as External Work
@@
From this perspective, retraining policies act as control strategies. Frequent retraining applies external work continuously, maintaining low mismatch at maximal operational cost, while infrequent retraining allows mismatch to accumulate and may permit performance to degrade beyond acceptable levels. Fixed schedules apply intervention independently of the underlying drift dynamics, whereas loss-triggered policies react only after degradation has already occurred.
Entropy-triggered retraining occupies an intermediate regime. By monitoring $D(t)$ (or a consistent estimator of it), retraining is initiated only when accumulated mismatch exceeds a tolerable threshold. This yields a principled tradeoff between predictive performance and retraining cost, applying external work selectively in response to the probability flow rather than to delayed performance signals.
@@ 5.1 Setup
We evaluate retraining strategies in a controlled binary classification task under synthetic nonstationarity. A logistic regression model is trained on an initial Gaussian feature distribution and deployed on a data stream with gradually shifting mean, increasing variance, and a rotating decision boundary.
We compare the following retraining strategies:
  • No retraining,
  • Daily retraining (maximum-frequency baseline),
  • Weekly retraining,
  • Performance-triggered retraining (label-based),
  • Entropy-triggered retraining (KL-based, label-free).
Retraining cost is measured by the number of retraining events, and predictive performance is measured by log loss.
@@ 5.2 Results
Figure 2 shows predictive log loss over time under drift. As expected, the model without retraining steadily degrades. Daily retraining provides a maximum-frequency baseline, keeping loss low at the cost of frequent retraining. Entropy-triggered retraining closely tracks the daily maximum-frequency baseline with far fewer retraining events, indicating that the mismatch signal provides actionable advance warning.
Table 1 summarizes the performance–cost tradeoff. In this experiment, entropy-triggered retraining achieves average log loss within about one percent of daily retraining while reducing retraining count from $199$ to $10$ (roughly a $95\%$ reduction). Performance-triggered retraining achieves near-maximum-frequency baseline loss but retrains far more frequently, illustrating the operational burden of label-dependent reactive policies.
Figure 3 visualizes cumulative retraining counts over time. Daily and performance-triggered approaches incur near-continuous retraining, whereas entropy-triggered retraining concentrates interventions at a small number of drift-driven times. Finally, Figure 4 provides a Pareto view: entropy-triggered retraining lies close to the efficient frontier, offering near-maximum-frequency baseline performance at substantially reduced cost.
Figure 2: Predictive log loss over time under nonstationary drift for different retraining strategies. Entropy-triggered retraining tracks daily retraining while requiring far fewer interventions.
Method Avg Log Loss Worst Log Loss Final Log Loss # Retrains
No retraining 0.7716 1.5351 1.5286 0
Daily retraining 0.5586 0.6251 0.5118 199
Weekly retraining 0.5595 0.6252 0.5119 28
Entropy-triggered 0.5626 0.6252 0.5193 10
Performance-triggered 0.5592 0.6221 0.5144 138
Table 1: Predictive performance and retraining cost under nonstationary drift.
Figure 3: Cumulative number of retraining events over time. Entropy-triggered retraining reduces retraining frequency by an order of magnitude relative to daily and performance-triggered baselines.
Figure 4: Cost–performance tradeoff (Pareto view). Each point corresponds to a policy, with cost measured by the number of retraining events and performance measured by average log loss.
@@ 7 Conclusion
We presented a theory-driven retraining framework that interprets model degradation under data drift as nonequilibrium entropy production. Modeling drift as Fokker–Planck probability flow yields a mismatch-entropy functional whose evolution is driven by probability currents and admits an entropy-balance decomposition with a nonnegative entropy production term. This provides a principled, label-free monitoring signal for deciding when retraining becomes necessary. Experiments demonstrate that entropy-triggered retraining achieves near-maximum-frequency baseline performance at a fraction of the retraining cost.
@@ Appendix A Proof of the Entropy Production Decomposition
In this appendix we derive the decomposition of the time derivative of the model mismatch entropy that underlies the entropy-triggered retraining framework. The structure of the result parallels entropy-balance relations in stochastic thermodynamics for diffusion processes. Related connections between entropy dissipation, diffusion, and variational structure are developed extensively in the optimal transport literature [Villani2003].
@@ Acknowledgements
The author used an artificial intelligence language model (ChatGPT, OpenAI; accessed December 29–31, 2025) as an assistive tool during manuscript preparation. The tool was used for expository editing and LaTeX formatting, limited symbolic checks of intermediate algebraic steps, and code drafting for experimental pipelines and figures. All mathematical derivations, modeling choices, experiments, and interpretations were independently derived, executed, and verified by the author.

### v2

@@ Abstract
Machine learning models deployed in nonstationary environments inevitably experience performance degradation due to data drift. While numerous drift detection heuristics exist, most lack a dynamical interpretation and provide limited guidance on how retraining decisions should be balanced against operational cost. In this work, we propose an entropy-based retraining framework grounded in nonequilibrium statistical physics. Interpreting drift as probability flow governed by a Fokker–Planck equation, we quantify model–data mismatch using relative entropy and show that its time derivative admits an entropy-balance decomposition featuring a nonnegative entropy production term driven by probability currents. Guided by this theory, we implement an entropy-triggered retraining policy using an exponentially weighted moving-average (EWMA) control statistic applied to a streaming kernel density estimator of the Kullback–Leibler divergence. We evaluate this approach across multiple nonstationary data streams. In synthetic, financial, and web-traffic domains, entropy-based retraining achieves predictive performance comparable to frequent retraining while reducing retraining frequency by one to two orders of magnitude. However, in a challenging biomedical ECG setting, the entropy-based trigger underperforms the maximum-frequency baseline, highlighting limitations of feature-space entropy monitoring under complex label-conditional drift.
@@ 1 Introduction
  • We empirically evaluate entropy-triggered retraining with an EWMA-based alarm on a streaming KL estimator, and quantify the resulting cost–performance tradeoffs across multiple drifting domains.
@@
Section 2 introduces a probability-flow model of drift. Section 3 defines mismatch entropy and derives its evolution along the flow. Section 4 interprets retraining as an external control and specifies the EWMA trigger used in experiments. Section 5 evaluates entropy-triggered retraining empirically. Appendix A provides the entropy-balance derivation, and Appendix B specifies the EWMA thresholding mechanism.
@@ 2 Drift as Probability Flow
Let $X_{t}\in\mathbb{R}^{d}$ denote the feature vector observed at deployment time $t$, where $t$ indexes a discrete-time data stream. In deployed systems, the underlying mechanisms generating drift are typically unknown and need not admit an explicit parametric description. Nevertheless, it is often useful to interpret distributional change through the lens of continuous-time stochastic dynamics.
As a conceptual model, we consider a stochastic process whose evolution is described by the stochastic differential equation
@@
where $a$ is a drift field, $B$ a diffusion matrix, and $W_{t}$ standard Brownian motion. This formulation is not assumed to govern the data-generating process in our experiments; rather, it serves as an interpretive framework for understanding how probability distributions evolve under persistent, nonequilibrium forcing.
Under standard regularity assumptions, the probability density $p(x,t)$ associated with such a process satisfies the Fokker–Planck equation
@@
The Fokker–Planck equation provides a canonical description of probability flow and admits a rich analytical structure; see [Risken1989] for a comprehensive treatment.
In deployed machine learning systems, neither the drift field $a$ nor the diffusion tensor $D$ is observed directly. Instead, the system is monitored through finite samples drawn from the time-varying feature distribution. From this perspective, the probability current $J(x,t)$ is not an observable object, but a conceptual quantity encoding the irreversible flow of mass in feature space induced by drift.
Nonzero probability currents correspond to nonequilibrium dynamics and provide a mechanism by which a fixed deployed model becomes misaligned with its environment. In nonequilibrium statistical physics, such currents are the fundamental drivers of irreversibility and entropy production, encoding sustained deviation from equilibrium through persistent probability flow [Schnakenberg1976]. This viewpoint motivates the use of entropy-based functionals as monitoring signals for deployment-time mismatch, even when only discrete-time empirical observations are available.
3 Mismatch Entropy and Its Dynamical Interpretation
@@
which measures the statistical divergence between the evolving deployment-time data distribution $p(x,t)$ and a fixed reference distribution $q_{\mathrm{ref}}$ associated with the deployed model (e.g., the training distribution or an updated reference after retraining). From an information-theoretic perspective, relative entropy provides a natural and asymmetric measure of distributional mismatch that is independent of labels or model outputs [CoverThomas2006].
Differentiating $D(t)$ along the Fokker–Planck probability flow yields
@@
More generally, along nonequilibrium probability flows the entropy rate admits a decomposition analogous to entropy-balance relations in stochastic thermodynamics, featuring a nonnegative entropy production term driven by probability currents and a residual “housekeeping” term that captures sustained external driving [Seifert2005]. In the present work, this decomposition serves as a theoretical motivation rather than a directly estimated quantity. The full derivation is provided in Appendix A.
@@
When $q_{\mathrm{ref}}$ coincides with an appropriate stationary reference for the flow, the mismatch entropy $D(t)$ behaves as a Lyapunov functional and decays monotonically. Under driven nonequilibrium drift, $D(t)$ need not be monotone; nevertheless, its evolution is constrained by an entropy-balance relation with a nonnegative entropy production term. Increases in mismatch reflect the effect of sustained probability currents relative to the fixed deployed reference distribution.
@@
However, it is important to note a fundamental limitation of this approach. The mismatch entropy $D(t)$ monitors changes in the marginal feature distribution $p(x)$ and is therefore most sensitive to covariate drift. In settings where performance degradation is driven primarily by changes in the conditional distribution $p(y\mid x)$ with relatively stable $p(x)$, feature-space entropy may fail to provide a timely or sufficient signal for retraining. This limitation is examined empirically in Section 5, where entropy-triggered retraining underperforms in a biomedical ECG domain characterized by strong label-conditional variability.
Figure 1: Evolution of an estimated mismatch entropy under drift. The threshold defines entropy-triggered retraining events. Small negative values may arise from finite-sample KDE estimation error; the true relative entropy is nonnegative.
@@
If the deployment-time distribution evolves under persistent probability currents, then mismatch relative to a fixed deployed reference will generally change over time, and sustained drift corresponds to nonequilibrium entropy production in the underlying dynamical system. Maintaining predictive performance therefore requires an intervention that updates the deployed reference. In deployed learning systems, this intervention takes the form of retraining.
4 Retraining as External Work and EWMA Triggers
@@
Practical KL estimator (as implemented)
In experiments, the mismatch entropy is estimated online using kernel density estimates (KDEs). At time index $t$ (discrete time), let $\{x_{i}\}_{i=1}^{n}$ denote a recent “fit” batch. We fit a KDE $\hat{p}_{t}$ on the fit batch and maintain a reference KDE $\hat{q}_{\mathrm{ref}}$ (updated only when an entropy-triggered retraining occurs). The empirical KL estimator used is
(6) $\widehat{D}_{t}=\frac{1}{n}\sum_{i=1}^{n}\Big(\log\hat{p}_{t}(x_{i})-\log\hat{q}_{\mathrm{ref}}(x_{i})\Big).$
EWMA thresholding
Instead of using a fixed sliding-window quantile threshold, retraining events are triggered using an exponentially weighted moving average (EWMA) alarm on the monitored statistic. Let $s_{t}$ denote a scalar monitoring signal (either $\widehat{D}_{t}$ for entropy-triggered retraining, or an online log-loss estimate for performance-triggered retraining). We maintain an EWMA mean $\mu_{t}$ and EWMA variance $v_{t}$:
(7) $\mu_{t}$ $=(1-\alpha)\mu_{t-1}+\alpha s_{t},$
(8) $v_{t}$ $=(1-\alpha)v_{t-1}+\alpha(s_{t}-\mu_{t})^{2}.$
The smoothing coefficient is parameterized by a half-life $h$ via
(9) $\alpha=1-2^{-1/h}.$
A retraining event is triggered when the standardized deviation exceeds a fixed level $k$:
(10) $z_{t}=\frac{s_{t}-\mu_{t}}{\sqrt{v_{t}}+\varepsilon}>k,$
with $\varepsilon>0$ a numerical stabilizer. In the experiments reported here, we use $h=50$ and $k=2.0$ for both entropy- and performance-based triggers, matching the experimental script.
@@ 5.1 Setup
We compare four retraining strategies implemented in a unified streaming evaluation pipeline:
  • No retraining: train once on an initial window and never update.
  • Daily retraining (maximum-frequency baseline): retrain at every time step on a rolling training window.
  • Entropy-triggered retraining: compute the KL estimator $\widehat{D}_{t}$ via KDE and trigger retraining using the EWMA rule in Section 4; on trigger, update both the classifier and the KDE reference.
  • Performance-triggered retraining: compute an online log-loss estimate and trigger retraining using the same EWMA rule; on trigger, update the classifier (labels assumed available).
Across all domains, the classifier is logistic regression trained on z-scored features. Feature normalization is fixed using the mean and standard deviation of the initial training window and applied to all subsequent data, including after retraining. Predictive performance is evaluated online using Bernoulli log loss, averaged over a small evaluation batch at each time step.
We evaluate these strategies on four drifting domains:
  • Synthetic Gaussian drift: two-dimensional Gaussian features with gradually shifting mean and variance, and a rotating logistic decision boundary. Each time step corresponds to a full block of samples.
  • Finance (SPY): daily market features (returns, rolling volatility, RSI, momentum) used to predict the next-day return sign over a long historical window.
  • wiki: daily pageviews for a popular article (Bitcoin), with log-transformed volume, rolling statistics, momentum, and weekly seasonality features.
  • Biomedical (MIT-BIH ECG): beat-level features extracted from ECG segments, labeled as normal versus abnormal beats using MIT-BIH annotations.
@@ 5.2 Results
Table 1 reports a cross-domain summary of average log loss and retraining frequency. Across the synthetic, financial, and web-traffic domains, entropy-triggered retraining achieves predictive performance close to the daily retraining baseline while requiring only a small fraction of the retraining events. In these settings, entropy monitoring successfully captures deployment-time drift that is relevant to predictive performance, allowing retraining to be applied selectively rather than continuously.
In contrast, on the MIT-BIH ECG domain, entropy-triggered retraining substantially underperforms daily retraining despite a large reduction in retraining frequency. This behavior reflects a structural limitation of feature-space entropy monitoring in this setting: clinically relevant changes in ECG data often manifest as shifts in the conditional distribution $p(x\mid y)$ rather than in the marginal feature distribution $p(x)$. High intra-class variability driven by patient-specific morphology, noise, and recording conditions can dominate the KDE-based KL estimate, reducing its sensitivity to decision-relevant drift. As a result, entropy-triggered retraining may respond too weakly or too late when predictive performance degrades, even though substantial label-conditional drift is present.
For clarity of presentation, time-series plots are shown for the Wikipedia Pageviews domain as a representative example. This domain exhibits smooth, interpretable covariate drift and a clear performance–cost tradeoff, making it well suited for visualizing the behavior of entropy-triggered retraining. Analogous qualitative behavior is observed in the synthetic and financial domains, whose quantitative results relative to the maximum-frequency baseline are reported in Table 1.
Domain Daily Avg Loss Entropy Avg Loss Entropy Retrains (%)
Synthetic Gaussian Drift 0.5587 0.5613 30 (15.0)
Finance (SPY) 0.6984 0.7016 90 (2.0)
Wikipedia Pageviews 0.6455 0.6377 68 (2.7)
Biomedical (MIT-BIH ECG) 0.1293 0.3224 76 (1.5)
Table 1: Cross-domain summary. “Daily” retrains at every step; percentages denote entropy-triggered retraining events relative to daily retraining within each domain.
Figure 2: Predictive log loss over time under drift for the Wikipedia Pageviews domain.
Figure 3: Cumulative number of retraining events over time for the Wikipedia Pageviews domain.
Figure 4: Cost–performance tradeoff (Pareto view) for the Wikipedia Pageviews domain.
@@ 6 Discussion
At the same time, the MIT-BIH ECG results indicate that a fixed, label-free drift signal and a fixed sensitivity $(h,k)$ can be insufficient for some domains. In such settings, more sensitive thresholds, domain-specific feature representations, or hybrid triggers combining supervised and unsupervised monitoring may be necessary.
@@ 7 Conclusion
We presented a theory-driven retraining framework that interprets model degradation under data drift as nonequilibrium entropy production. Modeling drift as Fokker–Planck probability flow yields a mismatch-entropy functional whose evolution is driven by probability currents and admits an entropy-balance decomposition with a nonnegative entropy production term. This provides a principled, label-free monitoring signal for deciding when retraining becomes necessary. Experiments demonstrate that entropy-triggered retraining with EWMA thresholding can achieve near-maximum-frequency baseline performance at a fraction of the retraining cost in several domains, but also reveal failure modes in which frequent retraining remains necessary.
@@ Appendix A Proof of the Entropy Production Decomposition
In this appendix we derive the decomposition of the time derivative of the model mismatch entropy that underlies the entropy-triggered retraining framework. The structure of the result parallels entropy-balance relations in stochastic thermodynamics for diffusion processes.
@@
Appendix B EWMA Thresholding Details
This appendix records the EWMA trigger used for both entropy- and performance-triggered retraining in the experimental script.
Let $h$ be the half-life and define $\alpha=1-2^{-1/h}$. Let $s_{t}$ be the monitored scalar statistic at time $t$ (either $\widehat{D}_{t}$ or an online log-loss estimate). The EWMA recursion is
(26) $\mu_{t}$ $=(1-\alpha)\mu_{t-1}+\alpha s_{t},$
(27) $v_{t}$ $=(1-\alpha)v_{t-1}+\alpha(s_{t}-\mu_{t})^{2}.$
We trigger retraining when $z_{t}>k$, where
(28) $z_{t}=\frac{s_{t}-\mu_{t}}{\sqrt{v_{t}}+\varepsilon}.$
In experiments, $h=50$, $k=2.0$, and $\varepsilon=10^{-12}$.
@@ Acknowledgements
The author used an artificial intelligence language model (ChatGPT, OpenAI; accessed December 29–31, 2025) as an assistive tool during manuscript preparation. The tool was used for expository editing and formatting, limited symbolic checks of intermediate algebraic steps, and code drafting for experimental pipelines and figures. All mathematical derivations, modeling choices, experiments, and interpretations were independently derived, executed, and verified by the author.